An airplane pilot finds he can go 72 miles in 35 minutes when traveling with the wind. He finds it takes him 44 minutes at the same airspeed to go 57 miles against the same wind. Find the speed of the wind and the airspeed of the plane.
Answer by checkley77(12844)

72/35=2.057 mi/m. with the wind.
57/44=1.295 mi/m. against the wind.
(2.057-1.295)/2=.762/2=.381 mi/m. is the speed of the wind
2.057-.381=1.676 mi/m. is the speed of the plane in still air.
1.295+.381=1.676 ditto.